17. Assume a bond is currently selling at par value. What will happen if the bonds expected cash flows are discounted at a rate lower than the bonds coupon rate? A. The price of the bond will increase B. The C. The par value of the bond will decrease. D. The coupon payments will be adjusted to the new discount coupon rate of the bond will increase. 18. As the discount rate is increased, the NPV of a specific project will A. increase. B. decrease. C. remain constant. D. decrease o zero, then remain 19. How much interest is earned in just the third year on a $1,000 deposit that earns 7% interest compounded annually? A. $70.00 B. $80.14 C. $105.62 D. $140.00 20. Which one of these costs accounts for the difference between accounting income and economic value added? B. Cost of capital C. Taxes D. Dividends 21. Net working capital is calculated by taking the difference between A. total assets and total liabilities. B. inventory and accounts payable C. current assets and curent liabilities D. cash and accounts payable. 22. After-tax operating income for a leveraged firm is defined as: A. net incme after-tax interest B. EBITx (1 tax rate) C. net income + depreciation. D. profit margin sales
Answer #1

17

The Price of the Bond will increase

Because bond coupon rate is more than the discount rate.

18

Decrease

Because of higher discount rate the NPV will be lower

19

Future value at year =1000*(1+7%)^2

Interest earned at year3 =1000*(1+7%)^2*7%=$80.14